# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 10.6.1 Smart Reserves and Meritorious Horizontal Choice Rule
We now introduce a generalization of the minimum guarantee choice rule through a smart reserve system. The key innovation is to allocate all HR-protected positions simultaneously by awarding units to the highest merit-score individuals, while respecting certain constraints. Specifically, these constraints involve (1) providing minimum guarantees for each HR-protected group and (2) maximizing the total number of HR-protected positions that are honored.
@@
Let’s illustrate this procedure using Example 8. We will consider individuals one at a time based on their merit scores, from highest to lowest. Under the minimum guarantee policy, HR-protected positions must be awarded before the remaining positions in the category.
@@
The highest merit-score individual, Amita, belongs to both HR-protected groups, so she must receive an HR-protected position, which could be from either group. Instead of immediately fixing her assignment to a specific position, we leave it flexible for now, determining which position she receives based on which assignment allows higher merit-score individuals to secure positions further down the merit ranking.
@@
The second-highest merit-score individual, Bhaskar, is not a member of any HR-protected group; he will have to wait, although he is certain to receive an open category position, as there is one untied to any HR-protected group. Next is Chandra, who belongs to only one HR-protected group: W. She is assigned the Open-W position, which means Amita must receive the Open-PwD position. With the HR-protected positions now settled, Bhaskar—the highest merit-score individual who has not yet received a position—receives the remaining open position.
@@
Critically, if the scenario were to be slightly different and Debraj—a man with a disability—had a higher merit score than Chandra, he would receive the Open-PwD position, resulting in Amita being assigned the Open-W position instead. Leveraging the flexibility in Amita’s assignment to allocate positions to higher-merit applicants is the “smart” aspect of this procedure.
@@
This idea extends naturally regardless of the specifics of the HR-protected groups or their overlapping memberships. First, we allocate the HR-protected positions in a “greedy” manner to the highest merit-score individuals, maximizing the total number of these positions that are honored.124124124Technically, this requires solving a maximum matching problem in a bipartite graph. In an unweighted bipartite graph, the objective is to find a matching with maximum cardinality. The Hopcroft-Karp algorithm (Hopcroft and Karp, 1973) solves this problem efficiently, and even faster algorithms exist for specific classes of bipartite graphs. Then, we allocate the remaining positions in the category to the highest merit-score eligible individuals who have not yet received a position. Introduced in Sönmez and Yenmez (2020), we refer to this choice rule as the meritorious horizontal choice rule.125125125The terminology meritorious horizontal choice rule was first introduced in Sönmez and Yenmez (2022). Originally, Sönmez and Yenmez (2020) referred to this choice rule as the horizontal envelope choice rule.
@@
Consider a setting with a single category and multiple HR-protected groups with overlapping memberships. Then, the meritorious horizontal choice rule is the only choice rule that satisfies non-wastefulness, no justified envy, and maximal accommodation of HR protections.
10.6.2 Two-Step Meritorious Horizontal Choice Rule
Having clarified the proper generalization of the minimum guarantee choice rule for the version of the problem with overlapping HR protections, a natural extension of the 2SMG mechanism emerges: the two-step meritorious horizontal (2SMH) choice rule.
Under this procedure, open category positions are allocated in the first step using the meritorious horizontal choice rule. In the second step, VR-protected positions are allocated, category by category, to eligible individuals who did not receive an open position in the first step, again using the meritorious horizontal choice rule.
@@ 10.6.2 Two-Step Meritorious Horizontal Choice Rule
Consider a setting in which individuals can belong to multiple HR-protected groups. Then, 2SMH is the only choice rule that satisfies non-wastefulness, no justified envy, maximal accommodation of HR protections, and compliance with VR protections.
By Theorem 14, and subject to the caveats discussed earlier in this section—namely, when objections to assignments may involve third parties—there exists a unique choice rule that fulfills the mandates of Saurav Yadav vs. State of U.P. (2020) even when the HR-protected groups are overlapping.
@@ 10.7 Allocation of Heterogenous Positions
Up to this point, our focus in Section 10 has been on problems involving identical positions. Another common variant in India, however, involves the centralized allocation of heterogeneous positions across multiple institutions. This variant is most prevalent in the allocation of prestigious government jobs—such as those in the Indian Administrative Service, Indian Police Service, and Indian Forest Service—and in the assignment of college seats at medical and engineering schools. As in the simpler case with identical positions, VR and HR policies are also enforced here, as codified in Indra Sawhney vs. Union of India (1992). Unlike the case of identical positions, however, the Supreme Court refrained from formulating and enforcing an explicit procedure to implement these policies for heterogeneous positions.
@@
The creation of such “artificial” property rights for specific categories of individuals based on flawed mechanics not only represents a misguided attempt to resolve the underlying issues with the SSD-based mechanism but also directly contradicts the axiom of no justified envy. Thus, the root cause of these legal challenges lies in the excessive reliance on the concept of “migration,” further exacerbated by the practice of awarding “artificial” property rights to arbitrary groups based on the tentative assignments of meritorious reserved candidates. This reliance on migration is an ill-equipped tool for addressing applications with heterogeneous positions and represents a methodological failure at the core of the legal inconsistencies and litigations in India.126126126A simple search on Indian Kanoon, a free search engine for Indian law, reveals that as of November 2024, there have been 959 cases related to the “migration of meritorious reserved candidates” at the Supreme Court or various high courts.
To address these challenges, it is important to recognize that accommodating the principles outlined in Indra Sawhney vs. Union of India (1992) or their refinements in Saurav Yadav vs. State of U.P. (2020) cannot be achieved solely by relying on the concept of migration—unless an arbitrary number of rounds of migrations are allowed through an iterative procedure such as the individual-proposing Deferred Acceptance algorithm. The necessity of iterative procedures is analogous to the need for such methods to find stable outcomes in two-sided matching markets.
10.7.1 A Minimalist Resolution: 2SMH-DA Mechanism
Even though Saurav Yadav vs. State of U.P. (2020) pertains to the version of the problem with identical positions, it also clarifies ambiguities in Indra Sawhney vs. Union of India (1992) concerning VR and HR policies, including for the more general version with heterogeneous positions. This observation suggests a minimalist resolution for this more general problem with heterogeneous positions and overlapping HR-protected groups: using the individual-proposing Deferred Acceptance algorithm (see Section 6.4), where each institution determines its set of best applicants with the 2SMH choice rule. Let’s refer to this preference revelation mechanism as 2SMH-DA.
In essence, this problem can be thought of as the more basic version of the problem involving identical positions, embedded within the school choice problem (see Section 6)—particularly the Turkish student placement variant, where the axiom of no justified envy plays an indispensable role. In this sense, 2SMH-DA, which simply integrates the respective solutions for these two applications, emerges as a very natural solution to India’s affirmative action problem with heterogeneous positions.
To further illustrate why this mechanism is uniquely appealing for this version of the problem, we extend the following two axioms from Saurav Yadav vs. State of U.P. (2020) into this more general setting.
@@ 10.7.1 A Minimalist Resolution: 2SMH-DA Mechanism
An allocation of positions across vertical categories (including open) at multiple institutions satisfies no justified envy if there exists no institution-category pair $(b,v)$ and two applicants $i$ and $j$, both eligible for category $v$ at institution $b$, such that:
@@
    Applicant $j$ is awarded a position at institution $b$ in category $v$,
@@
    Applicant $i$ strictly prefers a position at institution $b$ to their current assignment, and
@@
    Applicant $j$ neither has a merit score as high as that of applicant $i$, nor would awarding them a position at institution $b$ in category $v$ at the expense of applicant $i$ increase the number of HR-protected positions honored at institution $b$ in category $v$.
@@
An allocation of positions across vertical categories (including open) at multiple institutions satisfies maximal accommodation of HR protections if, for any individual $i$, institution $b$, and vertical category $v$, whenever assigning individual $i$ a position in category $v$ at institution $b$ (instead of their current assignment, possibly replacing another individual who was assigned a position in category $v$ at institution $b$) increases the number of HR-protected positions honored in that category at institution $b$, it must be the case that individual $i$ strictly prefers their current assignment over institution $b$.
Our next two results establish that, although multiple mechanisms satisfy the mandates of Saurav Yadav vs. State of U.P. (2020) for the more general version of the problem in India with heterogeneous positions, 2SMH-DA remains the only plausible option among them.
@@
Consider a setting with heterogeneous positions and overlapping HR-protected groups. Then, 2SMH-DA satisfies individual rationality, non-wastefulness, no justified envy, maximal accommodation of HR protections, and compliance with VR protections, and it Pareto dominates any other mechanism that also satisfies these axioms.
Theorem 16 (Sönmez and Yenmez, 2024).
Consider a setting with heterogeneous positions and overlapping HR-protected groups. Then, 2SMH-DA is the only mechanism that satisfies individual rationality, non-wastefulness, no justified envy, maximal accommodation of HR protections, compliance with VR protections, and strategy-proofness.
@@ 11 Medical Resource Rationing During COVID-19
The COVID-19 pandemic revealed another vital resource allocation problem, where minimalist market design proved highly effective: emergency rationing of scarce medical resources (Pathak et al., 2024). Thanks to the central focus this approach places on the objectives of the stakeholders, over the course of a year, we were able to communicate our ideas with dozens of healthcare officials when these ideas were most useful. The same focus also helped our team to convince many of these officials or experts in medical ethics and public healthcare to advocate for and adopt the reserve systems we designed for equitable allocation of vaccines and therapies on real time, as the need arise (Persad et al., 2020; Schmidt, 2020; Schmidt et al., 2020; Sönmez et al., 2021; Schmidt et al., 2021; Rubin et al., 2021; Pathak et al., 2021, 2022; White et al., 2022; McCreary et al., 2023).
@@ 11.1 The Toughest Triage
Upon exploring with Parag Pathak and Utku Ünver how these ethical principles were operationalized for allocation of ventilators and ICUs, we made an important discovery, one that made it clear that it is time to put our expertise into good use (Pathak et al., 2020b). Little did we know that, for nearly a whole year from that time on, the three of us would end up setting aside all unessential activities and devote all our time and energy to contribute to the global fight to combat COVID-19.
@@ 11.2 Priority Point System: Prevalence, Challenges, and Constraints
As their recommended allocation rule for crisis rationing of ventilators and ICUs, state guidelines in the U.S. almost uniformly suggest some form of a priority point system. In these guidelines, a patient’s point score is most often determined by a measure of mortality risk called the Sequential Organ Failure Assessment (SOFA) score.127127127This procedure can be thought as a simple serial dictatorship induced by the SOFA score. I refer to this specific priority point system as the SOFA system.128128128SOFA scoring system is first introduced in Vincent et al. (1996) on behalf of the Working Group on Sepsis-Related Problems of the European Society of Intensive Care Medicine. Under this system, each of six organ groups lungs, liver, brain, kidneys, blood clotting and blood pressure is assigned a score of 1 to 4, with higher scores for more severely failed organs. The total across the six organ groups determine the SOFA score of the patient, which then determines her priority for a ventilator or an ICU. In recent years, however, many expressed concerns that the SOFA system discriminates against certain groups. See, for example, Tolchin et al. (2021) for one such study.
@@
For example, at the onset of COVID-19, the leading allocation system that accommodated multiple ethical principles for the allocation of ventilators was a priority point system by White et al. (2009). Through a modification of the SOFA system, their system incorporated three ethical principles as follows: It utilized a coarsening of patient SOFA scores to accommodate the principle of saving most lives and added points to address two additional principles, saving the most years of life and the life-cycle principle.129129129The life-cycle principle asserts that younger patients, who have had the least opportunity to live through various stages of life, should be given higher priority for the scarce medical resource.
@@ 11.3 Pandemic Reserve System: A Minimalist Resolution
In real-world settings, reserve systems have been used for decades as a prominent tool for reaching compromises between groups with competing objectives. These settings include ensuring equitable representation of minorities and women in legislative assemblies (Htun, 2004; Dahlerup, 2007; Krook, 2009; Hughes, 2011), implementing affirmative action for allocating government positions and public school seats (Correa et al., 2019; Dur et al., 2020; Aygün and Bó, 2021; Sönmez and Yenmez, 2022), balancing school seat priorities between neighborhood populations and the broader community (Dur et al., 2018b), and allocating immigration visas in the U.S. (Pathak et al., 2025). However, given the diverse allocation criteria for different medical resources across numerous jurisdictions, the urgent need to allocate scarce medical resources globally during the pandemic required the examination of a more generalized class of reserve systems (Pathak et al., 2024).
@@
While the task forces in New York State and Minnesota could have adopted a similar approach by allocating a fraction of ventilators with heightened priority for essential personnel, a more flexible reserve system was generally needed. This flexibility was crucial for implementing the ethical principles outlined in Emanuel et al. (2020). Our new application of market design in public healthcare required a reserve system in which the priority ranking for any portion of the scarce resource could be determined by various, potentially unrelated criteria, without relying on a fixed baseline ranking.
Collaborating with Bumin Yenmez and working around the clock to make a timely contribution to the global fight against COVID-19, Pathak, Ünver, and I completed and circulated the first version of our paper Pathak et al. (2020c) in April 2020, just two weeks after beginning this critical resource allocation project. The final version was later published in Pathak et al. (2024).
@@
While the concept of a reserve system is intuitive, as we have previously discussed in depth, it contains a subtle aspect that can cause unintended distributional consequences if not properly understood by system operators: specifically, what happens when an individual has sufficiently high priority for the scarce resource across multiple categories?
@@
As a side note, while Pathak et al. (2020c) marked the first time we used minimalist market design in real-time to offer a resolution to an important crisis, it was not the last. When the contested 103rd Constitutional (Amendment) Act, 2019 was deliberated in the hearings of a recent Supreme Court judgment Janhit Abhiyan vs. Union of India (2022) in November 2022, in Sönmez and Ünver (2022), we offered a real-time resolution before a divisive decision was declared by the court in a split 3-2 vote, with strong protests from the two dissenting justices.130130130Indeed, about a week before the split decision, in Deshpande and Sönmez (2023), I collaborated with Ashwini Deshpande, a prominent expert on the Indian reservation system, to publish an op-ed in The Hindu. Together, we cautioned Indian society on the subtleties of the case. These experiences highlight the utility of minimalist market design—a highly structured framework based on identifying and remedying root causes of failures—in providing timely resolutions for important crises when they are most useful for society.

### v4

@@ 10.6.1 Smart Reserves and the Two-Step Meritorious Horizontal Choice Rule
The key innovation is to allocate all HR-protected positions simultaneously by awarding units to the highest merit-score individuals, while respecting certain constraints. Specifically, these constraints involve (1) providing minimum guarantees for each HR-protected group and (2) maximizing the total number of HR-protected positions that are honored.
@@
Let’s illustrate this procedure using Example 13, depicted in Figure 29. We will consider individuals one at a time based on their merit scores, from highest to lowest. Under the minimum guarantee policy, HR-protected positions must be awarded before the remaining positions in the category.
@@
The highest merit-score individual, Amita, belongs to both HR-protected groups, so she must receive an HR-protected position, which could be from either group. Instead of immediately fixing her assignment, we leave it flexible, determining which position she receives based on the allocation that enables higher-ranked individuals to secure positions later.
@@
The second-highest merit-score individual, Bhaskar, is not a member of any HR-protected group; he will have to wait, although he is certain to receive an open category position, as there is one not tied to any HR-protected group.
Next is Chandra, who belongs to only one HR-protected group: W. Here Amita’s assignment becomes decisive: for Chandra to receive an HR-protected position, namely the Open–W position, Amita must be assigned the Open–PwD position (contrast Panel (b) with Panel (c) in Figure 29).
With the HR-protected positions now settled, Bhaskar—the highest merit-score individual who has not yet received a position—receives the remaining open position.
@@
Critically, if the scenario were slightly different and Debraj—a man with a disability—had a higher merit score than Chandra, he would receive the Open–PwD position, resulting in Amita being assigned the Open–W position instead. Leveraging the flexibility in Amita’s assignment to allocate positions to higher-merit applicants is the “smart” aspect of this procedure.
Meritorious Horizontal Choice Rule.
@@
This idea extends naturally regardless of the specifics of the HR-protected groups or their overlapping memberships. First, HR-protected positions are allocated in a “greedy” manner to the highest merit-score individuals, maximizing the total number of these positions that are honored.138138138Technically, this requires solving a maximum matching problem in a bipartite graph. In an unweighted bipartite graph, the objective is to find a matching with maximum cardinality. The Hopcroft-Karp-Karzanov algorithm (Hopcroft and Karp, 1973; Karzanov, 1973) solves this problem efficiently, and even faster algorithms exist for specific classes of bipartite graphs. Next, the remaining positions in the category are allocated to the highest merit-score eligible individuals who have not yet received a position. Introduced in Sönmez and Yenmez (2020), we refer to this choice rule as the meritorious horizontal choice rule.139139139The terminology meritorious horizontal choice rule was first introduced in Sönmez and Yenmez (2022). Originally, Sönmez and Yenmez (2020) referred to this choice rule as the horizontal envelope choice rule.
@@
Consider a setting with a single category and multiple HR-protected groups with overlapping memberships. Then, the meritorious horizontal choice rule is the unique choice rule that satisfies non-wastefulness, no justified envy, and maximal accommodation of HR protections.
Having clarified the proper generalization of the minimum guarantee choice rule for the version of the problem with overlapping HR protections, a natural extension of the 2SMG mechanism emerges.
Two-step Meritorious Horizontal (2SMH) Choice Rule.
In the first step, open-category positions are allocated using the meritorious horizontal choice rule. In the second step, VR-protected positions are allocated, category by category, to eligible individuals who did not receive an open position in the first step, again using the meritorious horizontal choice rule.
@@
Consider a setting in which individuals can belong to multiple HR-protected groups. Then, 2SMH is the unique choice rule that satisfies non-wastefulness, no justified envy, maximal accommodation of HR protections, and compliance with VR protections.
By Theorem 14, and subject to the caveats discussed earlier in this section—namely, when objections to assignments may involve third parties—there exists a unique choice rule that fulfills the mandates of Saurav Yadav (2020) even when the HR-protected groups are overlapping.
@@ 10.7 Allocation of Heterogenous Positions
Up to this point, our focus in Section 10 has been on settings with identical positions. Another common variant in India involves the centralized allocation of heterogeneous positions across multiple institutions. This variant is most prevalent in the allocation of prestigious government jobs—such as those in the Indian Administrative Service, Indian Police Service, and Indian Forest Service—and in the assignment of college seats in medical and engineering schools. As in the simpler case with identical positions, VR and HR policies are also enforced here, codified in Indra Sawhney (1992). However, unlike in the case of identical positions, the Supreme Court refrained from formulating and enforcing an explicit procedure for implementing these policies in the context of heterogeneous positions.
@@
The creation of such “artificial” property rights for specific categories of individuals based on flawed mechanics not only represents a misguided attempt to resolve the underlying issues with the SSD-based mechanism but also directly contradicts the axiom of no justified envy. Thus, the root cause of these legal challenges lies in the excessive reliance on the concept of “migration,” further exacerbated by the practice of awarding “artificial” property rights to arbitrary groups based on the tentative assignments of meritorious reserved candidates. This reliance on migration is an ill-equipped tool for addressing applications with heterogeneous positions and represents a methodological failure at the core of the legal inconsistencies and litigations in India.140140140A simple search on Indian Kanoon, a free search engine for Indian law, reveals that as of November 2024, there have been 959 cases related to the “migration of meritorious reserved candidates” at the Supreme Court or various high courts.
To address these challenges, it is important to recognize that accommodating the principles outlined in Indra Sawhney (1992) or their refinements in Saurav Yadav (2020) cannot be achieved solely by relying on the concept of migration—unless an arbitrary number of rounds of migrations are allowed through an iterative procedure such as the individual-proposing deferred acceptance algorithm. The necessity of iterative procedures is analogous to the need for such methods to find stable outcomes in two-sided matching markets.
10.7.1 A Minimalist Resolution: 2SMH–DA Mechanism
Even though Saurav Yadav (2020) pertains to the version of the problem with identical positions, it also clarifies ambiguities in Indra Sawhney (1992) concerning VR and HR policies, including for the more general version with heterogeneous positions. This observation suggests a minimalist resolution for this more general problem with heterogeneous positions and overlapping HR-protected groups: using the individual-proposing deferred acceptance algorithm (see Section 6.4), where each institution determines its set of best applicants with the 2SMH choice rule. Let’s refer to this preference revelation mechanism as 2SMH–DA.
In essence, this problem can be thought of as the more basic version of the problem involving identical positions, embedded within the school choice problem (see Section 6)—particularly the Turkish student placement variant, where the axiom of no justified envy plays an indispensable role. In this sense, 2SMH–DA, which simply integrates the respective solutions for these two applications, emerges as a very natural solution to India’s affirmative action problem with heterogeneous positions.
To further illustrate why this mechanism is uniquely appealing for this version of the problem, we extend the following two axioms from Saurav Yadav (2020) into this more general setting.
@@ 10.7.1 A Minimalist Resolution: 2SMH–DA Mechanism
An allocation of positions across vertical categories (including open) at multiple institutions satisfies no justified envy if there exists no institution-category pair $(s,v)$ and two applicants $i$ and $j$, both eligible for category $v$ at institution $s$, such that:
@@
    Applicant $j$ is awarded a position at institution $s$ in category $v$,
@@
    Applicant $i$ strictly prefers a position at institution $s$ to their current assignment, and
@@
    Applicant $j$ neither has a merit score as high as that of applicant $i$, nor would awarding them a position at institution $s$ in category $v$ at the expense of applicant $i$ increase the number of HR-protected positions honored at institution $s$ in category $v$.
@@
An allocation of positions across vertical categories (including open) at multiple institutions satisfies maximal accommodation of HR protections if, for any individual $i$, institution $s$, and vertical category $v$, whenever assigning individual $i$ a position in category $v$ at institution $s$ (instead of their current assignment, possibly replacing another individual who was assigned a position in category $v$ at institution $s$) increases the number of HR-protected positions honored in that category at institution $s$, it must be the case that individual $i$ strictly prefers their current assignment over institution $s$.
Our next result establishes that, although multiple mechanisms satisfy the mandates of Saurav Yadav (2020) for the more general version of the problem in India with heterogeneous positions, 2SMH–DA remains the only plausible option among them.
@@
Consider a setting with heterogeneous positions and overlapping HR-protected groups. Then:
  1. 1.
    2SMH–DA satisfies individual rationality, non-wastefulness, no justified envy, maximal accommodation of HR protections, and compliance with VR protections, and Pareto dominates any other mechanism that also satisfies these axioms.
  2. 2.
    2SMH–DA is the unique mechanism that satisfies individual rationality, non-wastefulness, no justified envy, maximal accommodation of HR protections, compliance with VR protections, and strategy-proofness.
@@ 11 Medical Resource Rationing During COVID-19
The COVID-19 pandemic revealed another vital resource allocation problem, where minimalist market design proved highly effective: emergency rationing of scarce medical resources (Pathak et al., 2024). Thanks to the central focus this approach places on stakeholder objectives, over the course of a year, we were able to communicate our ideas with dozens of healthcare officials when these ideas were most useful. The same focus also helped our team to convince many of these officials or experts in medical ethics and public healthcare to advocate for and adopt the reserve systems we designed for equitable allocation of vaccines and therapies on real time, as the need arise (Persad et al., 2020; Schmidt, 2020; Schmidt et al., 2020; Sönmez et al., 2021; Schmidt et al., 2021; Rubin et al., 2021; Pathak et al., 2021, 2022; White et al., 2022; McCreary et al., 2023).
@@ 11.1 The Toughest Triage
Upon exploring with Parag Pathak and Utku Ünver how these ethical principles were operationalized for the allocation of ventilators and ICUs, we made an important discovery—-one that made it clear it was time to put our expertise to good use (Pathak et al., 2020b). Little did we know that, for nearly a whole year from that point on, the three of us would end up setting aside all inessential activities and devoting all our time and energy to contributing to the global fight against COVID-19.
@@ 11.2 Priority Point System: Prevalence, Challenges, and Constraints
As their recommended allocation rule for crisis rationing of ventilators and ICUs, state guidelines in the U.S. almost uniformly suggest some form of a priority point system. In these guidelines, a patient’s point score is most often determined by a measure of mortality risk called the Sequential Organ Failure Assessment (SOFA) score.141141141This procedure can be thought as a simple serial dictatorship induced by the SOFA score. I refer to this specific priority point system as the SOFA system.142142142SOFA scoring system is first introduced in Vincent et al. (1996) on behalf of the Working Group on Sepsis-Related Problems of the European Society of Intensive Care Medicine. Under this system, each of six organ groups lungs, liver, brain, kidneys, blood clotting and blood pressure is assigned a score of 1 to 4, with higher scores for more severely failed organs. The total across the six organ groups determine the SOFA score of the patient, which then determines her priority for a ventilator or an ICU. In recent years, however, many expressed concerns that the SOFA system discriminates against certain groups. See, for example, Tolchin et al. (2021) for one such study.
@@
For example, at the onset of COVID-19, the leading allocation system that accommodated multiple ethical principles for the allocation of ventilators was a priority point system by White et al. (2009). Through a modification of the SOFA system, their system incorporated three ethical principles: it utilized a coarsening of patient SOFA scores to accommodate the principle of saving most lives and added points to address two additional principles, saving the most years of life and the life-cycle principle.143143143The life-cycle principle asserts that younger patients, who have had the least opportunity to live through various stages of life, should be given higher priority for the scarce medical resource.
@@ 11.3 Pandemic Reserve System: A Minimalist Resolution
Reserve systems have been prevalent in real-world settings as a prominent tool for reaching compromises between groups with competing objectives. Such settings include ensuring equitable representation of minorities and women in legislative assemblies (Htun, 2004; Dahlerup, 2007; Krook, 2009; Hughes, 2011); implementing affirmative action in the allocation of government positions and public school seats (Correa et al., 2019; Dur et al., 2020; Aygün and Bó, 2021; Sönmez and Yenmez, 2022); balancing school seat priorities between neighborhood populations and the broader community (Dur et al., 2018b); and allocating immigration visas in the U.S. (Pathak et al., 2025). However, given the diversity of allocation criteria for medical resources across jurisdictions, the urgent need to allocate scarce medical resources globally during the pandemic called for examining a more general class of reserve systems (Pathak et al., 2024).
@@
While the task forces in New York State and Minnesota could have adopted a similar approach by allocating a fraction of ventilators with heightened priority for essential personnel, a more flexible reserve system was generally needed. This flexibility was crucial for implementing the ethical principles outlined in Emanuel et al. (2020). Our new application of market design in public healthcare called for a reserve system in which the priority rankings for various portions of the scarce resource could be determined by criteria that were possibly, but not necessarily, unrelated—without relying on a fixed baseline ranking.
Collaborating with Bumin Yenmez and working around the clock to make a timely contribution to the global fight against COVID-19, Pathak, Ünver, and I completed and circulated the first version of our paper in April 2020 (Pathak et al., 2020c), just two weeks after beginning this critical resource allocation project. The final version was later published in Pathak et al. (2024).
@@
While the concept of a reserve system is intuitive, as we have previously discussed in depth, it contains a subtle aspect that can cause unintended distributional consequences if not properly understood by system operators: specifically, what happens when an individual has sufficiently high priority for the scarce resource across multiple categories? (See Example 11 and Figure 25.)
@@
As a side note, while Pathak et al. (2020c) marked the first time we used minimalist market design in real-time to offer a resolution to an important crisis, it was not the last. When the contested 103rd Constitutional (Amendment) Act, 2019 was deliberated in the hearings of a recent Supreme Court judgment Janhit Abhiyan vs. Union of India (2022) in November 2022, in Sönmez and Ünver (2022), we offered a real-time resolution before a divisive decision was declared by the court in a split 3-2 vote, with strong protests from the two dissenting justices.144144144Indeed, about a week before the split decision, in Deshpande and Sönmez (2023), I collaborated with Ashwini Deshpande, a prominent expert on the Indian reservation system, to publish an op-ed in The Hindu. Together, we cautioned Indian society on the subtleties of the case. Though aspired market design projects typically differ from commissioned ones in the flexibility of their timeframe, these experiences highlight the promise of minimalist market design in providing timely resolutions to important crises when they are most needed by society.